Exploring Symmetric Change Decompositons

Kathy O’Hara

Abstract: I will explore the fascinating world of Symmetric Chain Decompositions

Introduction: One of the greatest milestones in Combinatorics was the discovery of an explicit Symmetric Change Lattice for the ”Coarse” Young’s lattive [KOH]

References

[KOH] Kathleen O’Hara, ...

Shalosh B. Ekhad, c/o D. Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: ShaloshBEkhad at gmail dot com .

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: DoronZeil at gmail dot com .